Consider the country of Agrovia, whose GDP is discussed in “A Numerical Example” on page 396. Construct a set of national accounts like that in Table 20-6 assuming that wheat costs $5 per bushel, there is no depreciation, wages are three-fourths of national output, indirect business taxes are used to finance 100 percent of government spending, and the balance of income goes as rent income to farmers.
1
Expert's answer
2021-09-13T12:03:33-0400

Gross Domestic Product is measured on the basis of these three methods. GDP is the value of goods and services produced in an economy.

A) Product Method

B) Income Method

C) Expenditure Method

By doing calculation by these methods we get Gross Domestic Product.

The country Agrovia, GDP (Gross Domestic Product) to construct a set on National Accounts, the cost of wheat $ 5 Busshel, there is no depreciation wages has to be considered; are three fourth of national output, and indirect business taxes are 100 percent of government spending and the balance of income goes s rent income to farmers.


Solution:

The country Agrovia, GDP(Gross Domestic Product) to construct a set on National Accounts, the wheat cost $ 5 Busshel, there is no depreciation wages are three fourth of national output, indirect business taxes are 100 percent of government spending and the balance of income goes s rent income to farmers

GDP= 87 of C+10 of G+6 of I-3

=100

Wheat cost= 100x5= 500

Three fourth of wages= 500×3/4=375

Depreciation=0

Interest=10 = As it is 100% of government purchases

Rent= 115=500-(375+10)

GDP = wages+ Interest+ rent

=375+10+115

=500